Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 35     Entry time: 16 Feb 2023, 14:04
Author: Frederik Wolff-Fabris 
Group: XO 
Subject: SA2 status 

Data for the Horizontal jitter at the HED Pop-In was observed during beam delivery on cw05 (03-05.Feb) and cw07 (14-16.Feb) (see attached pictures). 

We have a similar base jitter in both periods of beam delivery to HED, ranging from a minimum of 50-60um to around 80-100um in both operational weeks.

 

Attachment 1: HED_PopinX_cw05_23.png  101 kB  | Hide | Hide all
HED_PopinX_cw05_23.png
Attachment 2: HED_PopinX_cw07_23.png  96 kB  | Hide | Hide all
HED_PopinX_cw07_23.png
ELOG V3.1.4-7c3fd00